OpenStudy (anonymous):

Find the variation constant and an equation of variation where y varies inversely as x and y=4 when x=16

OpenStudy (a_clan):

y=k/x 4=k/16 k=64

OpenStudy (anonymous):

the equation of variation is y=??

OpenStudy (a_clan):

y is a variable which varies inversely as x. the equation of variation is y=k/x.

OpenStudy (anonymous):

OOOHHH ok